Mr.Smith had $3000 invested at 5% per year. How much must he invest at 6% in order to make his total annual income $300?
Answer by checkley71(8403)

3000*.05+X*.06=300
150+.06X=300
.06X=300-150
.06X=150
X-150/.06
X=$2500 MUST BE ADDED @ 6% TO GET $300 INTEREST PER YEAR.
PROOF
3000*.05+2500*.06=300
150+150=300
300=300